Grade 10Discuss with Askiitians Tutors

please guide me about the formula

satyam pandey
14 Years agoGrade 10

1 Answer

Aman  Bansal
14 Years ago

Dear Satyam,

to solve this question, you have to consider some cases.

case I : let d=0

a+b+c=24.

You have to distribute 24 identical objects into 3 groups.no of solutions to this = (24+3-1)C(3-1)=26C2

case II : let d=1

a+b+c = 20

 

You have to distribute 20 identical objects into 3 groups.no of solutions to this = (20+3-1)C(3-1)=22C2

case III : let d=2

a+b+c=16

You have to distribute 16 identical objects into 3 groups.no of solutions to this = (16+3-1)C(3-1)=18C2

case IV : let d=3

a+b+c=12

You have to distribute 12 identical objects into 3 groups.no of solutions to this = (12+3-1)C(3-1)=14C2

case V : let d=4

a+b+c=8

You have to distribute 8 identical objects into 3 groups.no of solutions to this = (8+3-1)C(3-1)=10C2

case VI : let d=5

a+b+c=4

You have to distribute 4 identical objects into 3 groups.no of solutions to this = (4+3-1)C(3-1)=6C2

case VII : let d=6

a+b+c=0

You have to distribute 0 identical objects into 3 groups.no of solutions to this = 2C2

Hence total number of solutions = 1 +6C2 + 10C2 + 14C2 + 18C2 + 22C2 + 26C2

                                                    = 861

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert